Formulating Financial Statements from Raw Data Following is selected financial information from Abercrombie& Fitch for its fiscal year ended January 30, 2016 ( thousands) .. $1,844,461 Cash from operating activities Noncash assets Cash from investing activities. .. . Cash, ending year . . . $ 309,941 (119,504) 520,708 (122,567) Cash from financing activities*.... . . . . . . . . .. . .1,137,317 Total expenses (other than cost Cash from financing activities includes the effects of foreign exchange rate fluctuations Required a. Prepare the income statement, the balance sheet, and the statement of cash flows for Abercrombie & b. Do the negative amounts for cash from investing activities and cash from financing activities concern c. Using the statements prepared for part a, compute the following ratios (for this part only, use the year- Fitch for the fiscal year ended January 30, 2016 us? Explain end balance instead of the average for assets and stockholders equity): i. Profit margin ii. Asset turnover i Return on assets iv. Return on equity
Answer #1
Income Statement for Abercrombie & Fitch (in $)
Revenue 3518680
Less Cost of Goods Sold 1361137
Less Other Expenses 2118984
Net Profit 38559
Balance Sheet of Abercrombie & Fitch ($)
Shareholders' Funds 1295722
Other Liabilities 1137317
Total 2433039
Non cash Assets 1844461
Cash Assets 588578
Total 2433039
Cash Flow Statement of Abercrombie & Fitch (&)
Cash Balance in the beginning 520708
Cash flow from Operating Activities 309941
Cash flow from Investing Activities -122567
Cash flow from Financing Activities -119504
Cash Flow at the end 588578

(b). The negative balance from investing activities may indicate that the company is spending on capital assets, such as property, plant & equipment or stocks etc. It signifies that the cash outflow on purchase of assets exceeds the inflow from sale of assets. Now, this may be due to the entity's expansion plans keeping in view the future growth prospects. Hence, negative cash from investing activities is not necessarily a sign of danger.

The negative balance of financing activities means that the payments in the form of interest, dividends, discount on issue of securities and redemption of capital on premium exceed the incomes from investments. Such a situation may be an indicator of trouble in operational and financial management of the concern and calls for a detailed analysis.

(c). Profit Margin = Gross profit/ Revenue

= (Revenue - COGS)/Revenue

= (3518680 - 1361137)/3518680

=2157543/3518680

=61.32%

Asset Turnover = Net Sales/Total Assets

=2157543/2433039

=88.68%

Return on Assets = Net profit/ Total Assets

= 38559/2433039

=0.0158

Return on equity = net profit/shareholders' equity

=38559/1295722

= 0.0298